# Mingkai Liao

### v1

@@ 6 Conclusion
We presented PACIFIER, a graph reinforcement learning framework for pacing opinion polarization dynamics via sequential network interventions. PACIFIER formulates the canonical ModerateInternal and ModerateExpressed problems as sequential decision-making tasks on networked opinion dynamics, enabling adaptive and scalable moderation without relying on handcrafted, model-specific heuristics. Within a unified framework, PACIFIER further supports multiple FJ-consistent settings, including continuous-valued internal opinions, biased assimilation dynamics, and node removal. Empirical results on real-world Twitter follow and retweet networks demonstrate that PACIFIER consistently matches or outperforms strong baselines across datasets and intervention tasks.
This manuscript is an arXiv preprint intended to present the core framework and initial empirical evidence. In future revisions, we plan to expand the experimental suite, strengthen the analysis of generalization and robustness, and include additional ablations and comparisons to further clarify the strengths and limitations of the proposed approach.

### v2

@@ 6 Conclusion
We proposed PACIFIER, to our knowledge the first unified graph-learning framework, and in particular the first unified graph reinforcement learning framework, for opinion polarization moderation. In contrast to prior work in this area, which is largely dominated by analytical and model-specific optimization under the Friedkin–Johnsen (FJ) model, PACIFIER reformulates the canonical ModerateInternal (MI) and ModerateExpressed (ME) problems as graph-based sequential decision-making tasks under a practical one-shot planning regime.
To support effective learning in topology-preserving moderation, we introduced two key representation mechanisms: temporal-aware node marking, which makes intervention history explicit and resolves history-induced state aliasing, and polarization-aware auxiliary global features, which summarize opinion–structure interaction patterns without requiring repeated steady-state recomputation. These components enable a shared encode–decode architecture that supports both long-horizon learning (PACIFIER-RL) and myopic ranking (PACIFIER-Greedy).
Experiments on polarized real-world Twitter follow and retweet graphs reveal a clear regime-dependent picture. In analytically structured MI settings, PACIFIER remains competitive with strong analytical solvers and consistently emerges as the strongest scalable non-analytical alternative. In contrast, in ME, continuous-ME, and cost-ME, PACIFIER shows strong and highly consistent superiority over non-PACIFIER baselines. Most importantly, PACIFIER-RL becomes decisively superior when long-horizon planning truly matters, most notably in cost-ME and topology-altering node_removal, where explicit reasoning over future structural consequences is crucial.
These results show that the contribution of PACIFIER is not only algorithmic but also conceptual: it helps identify when polarization moderation remains primarily an analytical optimization problem, and when it should instead be treated as a sequential, path-dependent planning problem. Future work includes extending PACIFIER to multi-camp and multi-topic settings, richer intervention primitives, and a deeper theoretical understanding of when learning can match or surpass analytical methods. Overall, PACIFIER shifts opinion polarization moderation from a domain dominated by model-specific analytical optimization toward a unified graph-learning and graph-reinforcement-learning paradigm.